OpenStudy (anonymous):

Two equations are shown: • Equation 1: 3/4(x−12)=12 •Equation 2: 3/4y−12=12 Solve each equation. Then, enter a number in each box to make this statement true. The value of x is: The value of y is:

geerky42 (geerky42):

\[\dfrac34(x-12)=12\] We need to isolate x. Any idea what to do with \(\dfrac34\)?

OpenStudy (anonymous):

convert it to a decimal?

geerky42 (geerky42):

Sure. Then we have \(.75(x-12) = 12\) We still need to get rid of .75

geerky42 (geerky42):

Maybe do something to both sides to cancel .75 out?

OpenStudy (anonymous):

subtract .75 from .75 and 12?

geerky42 (geerky42):

I would go with dividing.

geerky42 (geerky42):

\(\dfrac{.75(x-12)}{.75} = \dfrac{12}{.75}\) \(\dfrac{\cancel{.75}(x-12)}{\cancel{.75}} = \dfrac{12}{.75}\) \(x-12 = \dfrac{12}{.75}\)

OpenStudy (anonymous):

okk

geerky42 (geerky42):

Do you understand what is going on?

OpenStudy (anonymous):

yes so 12 divided by .75 is 16?

OpenStudy (anonymous):

so then 12/.75 is equal; to 9, 

OpenStudy (anonymous):

okk so x=28? 28-12=16 witch is 16=16? right

OpenStudy (anonymous):

yes that right :)

OpenStudy (anonymous):

yay thank u guys!!!!! :-)

OpenStudy (anonymous):

now equation 2 I need help also

OpenStudy (anonymous):

do I do the same thing as I did the first time?

OpenStudy (anonymous):

no its then 3/4Y-12 = 12 you add 12 to both sides and its 3/4Y = 24 you divide btoh sides by 3/4 and its Y = 32

OpenStudy (anonymous):

its a lot to chew so if you get stuck ill help you with each part

OpenStudy (anonymous):

yea I'm a little confused can u explain please

OpenStudy (anonymous):